OpenStudy (caitlinnr14):

WILL GIV MEDALLLLL! If h(x) = −3x − 10, find h(−3). −19 −1 1 2.3 I got -1

OpenStudy (anonymous):

Your are correct, it is -1

OpenStudy (anonymous):

yes I THINK you're correct...

OpenStudy (anonymous):

replace the x's with -3, -3*-3=9 9-10=-1

OpenStudy (caitlinnr14):

ok THANK you @PFEH.1999

OpenStudy (caitlinnr14):

lol

OpenStudy (anonymous):

You re right !:)

OpenStudy (anonymous):

h(-3) = −3(-3) − 10 h(-3) = 9-10 h(-3)=-1

OpenStudy (caitlinnr14):

yay @E.ali

OpenStudy (caitlinnr14):

Thxx @bbbbbbbbbbbbbbbbbb

OpenStudy (anonymous):

Your welcome